OpenStudy (anonymous):

A clock starts at 5:05. What time is it after the minute hand rotates 12 degrees?

OpenStudy (anonymous):

simply view your clock as a circle a circle is 360˚ your clock has 60 minutes and hence is divided into 60 parts 360/60=6 the circle is 6 times 'bigger' in measurement than the clock so if you were to put the angle in a clock, it would look 6 times smaller let x be the number of minutes hence 12˚/360˚ = 6x/360 12=6x x=2 its 5:05 + 2 minutes

OpenStudy (anonymous):

Only in 3rd grade. Don't understand. Divide 360? Is there an easier way?

OpenStudy (jdoe0001):

hmmm how many degrees are there in a circle?

OpenStudy (anonymous):

360

OpenStudy (jdoe0001):

the minute hand goes around the clock/circle once and that's 1 hour and it moves a little bit well, in 1hr, there are 60minutes that means that in one way around the clock the minute hand has done 1hr, or 60minutes or 360 degrees so to know how many degrees the minute hand moves over 1 minute you'd grab the circle's 360 degrees and divide it by 60 even pieces

OpenStudy (jdoe0001):

\(\bf 360^o\div 60min\implies 6^o\ per \ min\)

OpenStudy (anonymous):

The question is: A clock starts at 5:05. What time is it after the minute hand rotates 12 degrees? Would it be

OpenStudy (anonymous):

This is the question- a clock starts at 5:05. What time is it after the minute hand rotates 12 degrees? Don't get it.

OpenStudy (anonymous):

if it rotates 6 degrees, its 1 minute therefore for 12 degrees, its 12/6=2 1 minute x 2 = 2minutes no idea why your school would give you such a question if you haven't learnt to divide.